# BPINN-EM-Post: Stochastic Electromigration Damage Analysis in the Post-Void Phase based on Bayesian Physics-Informed Neural Network

### v1

@@ IV-B Generating observations using analytical solutions
$\zeta_{1}(p,x,L)=(2p+2)L-x,\;\zeta_{2}(p,x,L)=(2p+1)L-x,$ (8)
$\zeta_{3}(p,x,L)=(2p)L+x,\;\;\;\;\;\;\;\zeta_{4}(p,x,L)=(2p+1)L+x,$
$g(x,t)=2\sqrt{\frac{kt}{\pi}}e^{-\frac{x^{2}}{4kt}}-x\times erfc\{\frac{x}{2\sqrt{kt}}\}$
For the segments with void, we assume that one end of the segment could either be an inter-segment junction or a terminal boundary, while the other end corresponds to a void node. Fig 2(b) illustrates a single segment with a void at the $x=L$ end. At $x=0$, which may represent either an inter-segment junction or a terminal boundary , a modifiable boundary condition from the voidless scenario is applied. We modify the boundary condition at $x=L$ to $\tilde{\sigma}(L,t)=0$. When using a very small $\delta$ in (LABEL:eq:korhonen_equation_single_wire_postvoid), the boundary condition at the site of the void mirrors the approach used here [41]. The boundary conditions for the segment with a void at the subsequent node of a wire segment, i.e., at $x=L$, are expressed as outlined in  (9).
$BC:\frac{\partial\tilde{\sigma}(x,t)}{\partial x}=\phi^{-}(t),\,\,x=0$ (9)
$BC:\tilde{\sigma}(x,t)=0,\,\,x=L$
Like the solution for the segment without a void, the solution for the single segment with a void at $x=L$ is derived using the Laplace transform method:
$\tilde{\sigma}(x,t)=$ (10)
$\sum_{p=0}^{P}(-1)^{p}\{\frac{d\phi^{-}(t)}{dt}*\left(g(\zeta_{1}(p,x,L),t)-g(\zeta_{3}(p,x,L),t)\right)$
$+\frac{2}{L}\sum_{m=1}^{M}(-1)^{m+1}\{e^{-M}\cos((m-0.5)\pi x/L)$
×$\times\int_{0}^{L}h(u)\sin\left((m-0.5)\pi(L-u)/L\right)du\}$
Similarly, when the void is positioned at the preceding (left/bottom) end, i.e., at $x=0$, the boundary condition is defined as $\tilde{\sigma}(x,t)=0$ at $x=0$ and $\frac{\partial\tilde{\sigma}(x,t)}{\partial x}=\phi^{+}(t)$ at $x=L$. Using the same Laplace transform method, the analytical solution for this scenario can also be derived.
In Eq (7) and Eq (10) the values $\phi^{-}(0)$ and $\phi^{+}(0)$ can be calculated using the current density information at each inter-segment junctions [28, 27]. For each single segments, we use the flux information from Bayesian network as $\frac{d\phi^{-}(t)}{dt}=F^{-}(\mathbf{x};\omega)$ and $\frac{d\phi^{+}(t)}{dt}=F^{+}(\mathbf{x};\omega)$.
The analytical solutions in Eq (7) and Eq (10) estimate the stress distribution for single segments within a multisegment interconnect structure, given the flux information. For accurate stress distribution, all relevant physical conditions must be satisfied. While calculating the flux information, atomic flux conservation was enforced. Additionally, as per Eq (1), stress continuity must be maintained at inter-segment junctions. This implies that the stress in all segments connected to a junction should be equal.
To ensure stress continuity, we optimize the BPINN using the stress continuity condition $\mathcal{L}$, defined as:
$\mathcal{L}=\frac{1}{N_{I}\times K_{i}}\sum_{i=1}^{N_{I}}\sum_{k=2}^{K_{i}}\left(\tilde{\sigma}_{k}(x_{\sigma};\omega)-\tilde{\sigma}_{k-1}(x_{\sigma};\omega)\right)^{2},$ (11)
where $\tilde{\sigma}_{k}(\omega)$ represents the electromigration (EM) stress at the boundary of the $k^{th}$ segment at inter-segment junction $i$, $K_{i}$ is the number of segments connected to junction $i$, and $N_{I}$ is the total number of inter-segment junctions in the multisegment structure. Since $\tilde{\sigma}$ depends on $F^{-}(\mathbf{x};\omega)$ and $F^{+}(\mathbf{x};\omega)$, it is expressed as a function of $\omega$ and the parameters influencing stress, denoted as $x_{\sigma}$, for simplicity.
IV-C posterior for BPINN
In this work, the Bayesian network takes time and current information at each inter-segment junction of the multisegment interconnect structure as input, i.e., $\mathbf{x}=\{T,J_{L},J_{U},J_{R},J_{D}\}$. Using the flux information from the network’s output, we enforce atomic flux conservation to calculate the flux values $F^{-}(\mathbf{x};\omega)$ and $F^{+}(\mathbf{x};\omega)$ at the endpoints of each segment.
Therefore for this case we can define the dataset as:
Here we assume that the observations are variational, and can be represented as:
¯$\bar{\mathcal{L}}=\mathcal{L}(x_{\mathcal{L}}^{(i)})+\epsilon^{(i)},i=1,2,3,..N^{\mathcal{L}}$ (13)
where $\epsilon$ is independent Guassian noise with zero mean and variance of $\mathbf{Var_{\mathcal{L}}}$. The likelihood of the observations can be calculated as:
$P(\mathcal{D}|\omega)=$ $\prod_{i=1}^{N_{\mathcal{L}}}\frac{1}{\sqrt{2\pi}Var_{\mathcal{L}}^{(i)}}$ (14)
×$\times exp\left[-\frac{\left(\mathcal{L}\left(x_{\mathcal{L}}^{(i)};w\right)-\bar{\mathcal{L}}^{(i)}\right)^{2}}{2Var_{\mathcal{L}}^{(i)}}\right]$
Utilizing this likelihood, the posterior can be obtained using Baey’s theorem as:
Here, the symbol $\propto$ signifies ’equality up to a constant’. The probability of the dataset $P(D)$ is not readily solvable through analytical methods. Therefore, in practical scenarios, we only obtain an un-normalized expression $P(\omega|\mathcal{D})$.
This work BPINN-EM-Post employs the Hamiltonian Monte Carlo (HMC) method to navigate the parameter space of our Bayesian Neural Network (BNN) model. HMC is a gradient-based Markov Chain Monte Carlo (MCMC) technique that utilizes Hamiltonian dynamics for parameter exploration. Our HMC procedure begins by simulating Hamiltonian dynamics through a numerical integration method, followed by a correction using the Metropolis-Hastings acceptance step. Given a dataset $\mathcal{D}$, we postulate the target posterior distribution for $\omega$ as follows:
$P(\omega|\mathcal{D})\simeq exp(-U(\omega))$ (16)
Here, $U(\omega)=-\log(P(\mathcal{D}|\omega))-\log(P(\omega))$ serves as the potential energy in the system. To sample from the posterior distribution, HMC introduces an auxiliary momentum variable $r$, which is used to formulate a Hamiltonian system:
$H(\omega,r)=U(\omega)+\frac{1}{2}r^{T}\mathrm{M}^{-1}r$ (17)
Here, $M$ is a mass matrix usually set to be the identity matrix, $I$. HMC then generates samples from the joint distribution of $(\omega,r)$ as:
$\pi(\omega,r)\sim exp(-U(\omega)-\frac{1}{2}r^{T}\mathrm{M}^{-1}r)$ (18)
Here, the samples of $\omega$ have a marginal distribution as we discard the samples of $r$. These samples are derived from the following Hamiltonian dynamics:
$d\omega=\mathrm{M}^{-1}rdt,$ (19)
$dr=-\nabla U(\omega)dt$
Eq (19) is discretized through the leapfrog method, and the Metropolis-Hastings step is employed to minimize the discretization error. More details on the HMC algorithm is explained in  [40].
IV-D Estimation of variations in EM stress
Using HMC, we sample the optimal Bayesian network parameters, $\omega$, to generate samples of flux information, $\left\{\tilde{\mathbf{F}}(\mathbf{x};\omega^{(i)})\right\}_{i=1}^{M}$. From these, we derive samples of $\{\mathbf{F}^{-}\}$ and $\{\mathbf{F}^{+}\}$. For each wire segment, these flux samples are used to compute $M$ stress solution samples, $\{\tilde{\sigma}(x,t)\}$.
Notably, when obtaining the optimal parameter samples via HMC, we utilize only stress observations at the boundaries of single wire segments within the multisegment interconnect structure. This significantly reduces the variables in the loss function $\mathcal{L}$, enabling efficient training. Once the framework is fully trained and optimal parameters are obtained, we calculate the flux information at all wire segment boundaries. This flux information is termed optimal as it satisfies flux conservation and stress continuity at inter-segment junctions. Using the optimal flux information samples, we compute accurate stress distribution samples for individual segments through analytical solutions. Additionally, variations in stress solutions, such as mean and standard deviation, are estimated from the calculated stress samples.
V Experimental results and discussion
(a) Stress distribution on single segment without void
(b) Stress distribution on a single segment with void at left node
(c) Stress distribution on a single segment with void at right node
Figure 3: Comparison of the solution obtained from analytical solution with the solution from FEM-based COMSOL for single interconnect segment for three different cases.
(a) Representation of a ten-segment multi-segment interconnect structure. The depicted current densities are the mean values of the current densities used during simulation.
(b) Comparison of variations estimation between the proposed method, COMSOL, and EMSpice at node 1 (node with void) of the depicted ten-segment interconnect.
(c) Comparison of variations estimation between the proposed method, COMSOL, and EMSpice at inter-segment junction 2 of the depicted ten-segment interconnect.
(d) Comparison of variations estimation between the proposed method, COMSOL, and EMSpice at inter-segment junction 5 of the depicted ten-segment interconnect.
(e) Comparison of variations estimation between the proposed method, COMSOL, and EMSpice at inter-segment junction 7 of the depicted five-segment interconnect.
Figure 4: Illustration of stress variations at internal junctions of a multi-segment interconnect.
The modules used in BPINN-EM-Post, such as the analytical solution, and the BPINN framework are developed using Python 3.9.18 with PyTorch 2.2.1. The training and testing procedures are conducted on a Linux server equipped with two Intel 22-core E5-2699 CPUs, 320 GB of memory, and an Nvidia TITAN RTX GPU.
V-A Data preparation and scaling
To determine the stress distribution in a multisegment interconnect structure, we first gather the physical and material attributes of the interconnect segments. Each segment $i$ is characterized by its physical properties, including length $L_{i}$ and width $w_{i}$, while assuming uniform thermal c onductivity $\kappa$ across all segments. Spatial points $x$ are sampled within $[0,L]$, and temporal points are chosen from $t=0$ to $1\times 10^{9}$ s at intervals of $1\times 10^{7}$ s. For this analysis, we assume that nucleation-phase solutions for these interconnects are pre-established by other methods, providing initial stress data $h(x)$ and void locations.
@@ V-A Data preparation and scaling
The input parameters for the analytical methods and the MLP vary significantly in scale, necessitating data scaling for standardization. Scaling constants $k_{x}$, $k_{t}$, and $k_{\sigma}$ are used for spatial ($x$), temporal ($t$), and stress ($\sigma$) parameters, respectively. Additional parameters are scaled using the following equation:
@@
$=$ $\sigma_{sc}(k_{x}x,k_{t}t,\frac{k_{x}^{2}}{k_{t}}\kappa,\frac{k_{\sigma}}{k_{x}}G),$
@@
To assess the accuracy and performance of BPINN-EM-Post, we compare results against Monte Carlo simulations using FEM-based COMSOL [39] and FDM-based EMSpice [18]. For each Monte Carlo iteration, branch currents are randomly sampled from a normal distribution with a mean range of $5\times 10^{-8}\,\mathrm{A/m^{2}}$ to $5\times 10^{9}\,\mathrm{A/m^{2}}$ to introduce variability in input currents.
V-B Accuracy and performance of analytical solutions
@@
In order to evaluate the precision and speed of the analytical solutions, we conducted comparisons with results from COMSOL simulations using a variety of specific stress gradients. We generated approximately 500 wire segments, with lengths varying from $10\mu m$ to $50\mu m$, all maintaining a uniform width of $1\mu m$. These segments were subjected to current densities spanning from $-5\times 10^{9}A/m^{2}$ to $5\times 10^{10}A/m^{2}$. We also set other parameters, including $e=1.6\times 10^{-19}C$, $Z^{*}=10$, $E_{a}=1.1eV$, $B=1\times 10^{11}$, $D_{0}=5.2\times 10^{-5}m^{2}/s$, $\rho=2.2\times 10^{-8}\Omega m$, $\Omega=8.78\times 10^{-30}m^{3}$, and $\sigma_{crit}=5\times 10^{8}Pa$ In the evaluation of our analytical solutions, the wire segments were modeled with non-zero boundary flux to represent their placement within the multisegment trees. For each segment, we collected data at 30 spatial positions $x\in[0,L]$ and across 100 time intervals from 0 to $1\times 10^{8}$ seconds to account for aging time $t$. These segments were individually analyzed using COMSOL and benchmarked against our analytical solution. We employed the root mean square error (RMSE) to measure discrepancies, noting that the stress values $\sigma(x,t)$ ranged from $-1.1\times 10^{8}$ to $4.6\times 10^{8}$. With our analytical approach, we achieved an average RMSE of $3.3\times 10^{4}$, correlating to an average error margin of 0.006%. The analytical solver recorded an average processing time of 0.01 seconds per segment, in stark contrast to COMSOL’s 7.5 seconds, resulting in a computational speedup of approximately 750 times. This demonstrates that our first-stage analytical solver delivers both rapid and precise results when juxtaposed with the finite element method (FEM)-based COMSOL simulations. Figure 3 illustrates the comparative results of our analytical solution versus COMSOL for segments both with and without voids, achieved using $M=5$ in the analytical calculations. This optimal value of $M=5$ was established through experimental determination and was consistently applied in all aspects of this study.
V-C Overall accuracy and performance analysis
@@
To assess the efficacy and precision of of our method across a variety of multi-segment interconnect structures, we analyzed 1000 such configurations, with segment counts varying between 50 and 250. These structures were examined to gauge variational EM stress through the use of COMSOL, EMSpice, and BPINN-EM-Post, utilizing 30 samples to measure variations in stochastic EM stress. The findings, detailing accuracy and performance metrics for multi-segment interconnect structures across different segment counts, are showcased in Table I. The observed range of EM stress across these structures spanned from -4.31×$\times 10^{9}$ to 5.94×$\times 10^{9}$. Accuracy was quantified using RMSE, computed as $RMSE=(RMSE_{mean}+RMSE_{std})/2$. For the dataset comprising 1000 multi-segment interconnect structures, the average RMSE relative to COMSOL stood at 7.07×$\times 10^{7}$, equating to an error margin of approximately 0.69%. In comparison, against EMSpice, the average RMSE was recorded at 9.5×$\times 10^{6}$, reflecting an error rate of about 0.093%.
TABLE I: Performance and accuracy comparisons with existing methods
# of wires COMSOL EMspice BPINN-EM-Post
Runtime (s) Runtime (s) Error(%) vs. Comsol Error(%) vs. EMSpice Inference time(s) Training time(s) Speedup vs. COMSOL Speedup vs. EMSpice
50 2560 876 0.25 0.051 0.52 6.5 365x 125x
100 4800 1490 0.31 0.074 0.71 15.3 300x 93x
150 6731 2122 0.52 0.094 0.91 21.8 296x 91x
200 9230 3100 0.74 0.126 1.25 37.9 236x 78x
250 10191 3656 1.1 0.163 1.51 55.6 178x 64x
@@
Figure 4 demonstrates the variations in EM stress, expressed through both mean and standard deviation, for a ten-segment interconnect structure, the physical and electrical characteristics of which are shown in Figure 4(a). The variations in EM stress at the selected junctions between segments of this interconnect structure are detailed in Figures 4(b), 4(c), 4(d), and 4(e) . In these comparisons, BPINN-EM is shown to precisely predict the variations in EM stress, aligning closely with the results from COMSOL and EMSpice.
@@
In the analysis of stochastic EM stress variances for multi-segment interconnect structures, Monte Carlo simulations are conducted for 30 cycles using both COMSOL and EMSpice. COMSOL’s average runtime for 1000 multi-segment interconnect structure is approximately 7500 seconds, while EMSpice takes about 2100 seconds in average. For BPINN-EM-Post, variations in stochastic EM stress are estimated by sampling from the Bayesian Neural Network (BNN) via Hamiltonian Monte Carlo (HMC). Prior to this, the analytical solutions must ascertain the optimal atomic flux at the termini of each segment within the structures, leveraging a Physics-Informed Neural Network (PINN). Thus, both the sampling(training) period of the PINN and the inference duration of the BNN are accounted for in the runtime evaluation. For the multi-segment interconnects assessed, the average runtime for BPINN-EM-Post is noted to be approximately 31 seconds, significantly accelerating the estimation process by a factor of 240 times faster than COMSOL and 67 times quicker than EMSpice. This efficiency marks our approach as vastly superior in terms of speed for calculating variational EM stress.
VI Conclusion
@@
In this paper, we introduced a novel machine learning-based stochastic analysis framework, termed BPINN-EM-Post, which integrates analytical solutions with a Bayesian Physics-Informed Neural Network (BPINN) to estimate variational Electromigration (EM)-induced stress effects. The framework combines closed-form analytical solutions with a physics-informed neural network to ensure stress continuity and atomic flux conservation during post-void stress calculations. BPINN-EM-Post effectively utilizes analytical solutions for individual wire segments, avoiding the need to model complex initial stress distributions. At the inter-segment level, physics constraints, such as stress continuity and atomic flux conservation, are enforced through the neural network, resulting in a more efficient and accelerated training process by reducing the number of variables in the loss function. The BPINN framework further enhances the approach by providing uncertainty quantification and enabling variational analysis, ensuring robust and efficient modeling of EM stress evolution. Experimental results demonstrate substantial computational improvements, achieving a 240×240$\times$240 × speedup compared to FEM-based Monte Carlo simulations and a 65×65$\times$65 × speedup over FDM-based methods. These findings underscore the potential of the BPINN framework for scalable and accurate variational EM stress analysis.

### v2

@@ IV-B Data Preparation and Scaling
[…]
@@
The input parameters for the analytical method and the MLP vary significantly in scale, necessitating data scaling for standardization. Scaling factors $k_{x}$, $k_{t}$, and $k_{\sigma}$ are used for spatial ($x$), temporal ($t$) and stress ($\sigma$), respectively. Additional parameters are processed in Eq. (20),
@@
$=\sigma_{sc}(k_{x}x,k_{t}t,\frac{k_{x}^{2}}{k_{t}}\kappa,\frac{k_{\sigma}}{k_{x}}G)$
@@
To assess the accuracy and performance of our proposed BPINN-EM-Post, we compare results against Monte Carlo simulations using FEM-based COMSOL [16] and FDM-based EMSpice tool [13]. For each Monte Carlo iteration, branch currents are randomly sampled from a normal distribution with a mean range of $5\times 10^{-8}\,\mathrm{A/m^{2}}$ to $5\times 10^{9}\,\mathrm{A/m^{2}}$ to introduce variability to input currents.
IV-C Accuracy and Performance of Analytical Solutions
@@
In order to evaluate the precision and speed of the analytical solutions, we conducted comparisons with results from COMSOL simulations using a variety of specific stress gradients. We randomly generated 500 single wire segments samples, with lengths varying from $10\mu m$ to $50\mu m$, all maintaining a uniform width of $1\mu m$. These segments were subjected to current densities spanning from $5\times 10^{9}A/m^{2}$ to $5\times 10^{10}A/m^{2}$. Other physical parameters are set as $e=1.6\times 10^{-19}C$, $Z^{*}=10$, $E_{a}=1.1eV$, $B=1\times 10^{11}$, $D_{0}=5.2\times 10^{-5}m^{2}/s$, $\rho=2.2\times 10^{-8}\Omega m$, $\Omega=8.78\times 10^{-30}m^{3}$, and $\sigma_{crit}=5\times 10^{8}Pa$. In the evaluation of our analytical solutions, the wire segments were modeled with non-zero boundary flux to represent their placement within the multi-segment trees. For each segment, we collected data at 30 spatial positions $x\in[0,L]$ and across 100 time intervals from 0 to $1\times 10^{8}$ seconds to account for aging time $t$. These segments were individually analyzed using COMSOL and benchmarked against our analytical solution. We employed the root mean square error (RMSE) to measure discrepancies, noting that the stress values $\sigma(x,t)$ ranged from $1.1\times 10^{8}$ to $4.6\times 10^{8}$ Pa. With our analytical approach, we achieved an average RMSE of $\mathbf{3.3\times 10^{4}}$ Pa, correlating to an average error margin of 0.006%. The analytical solver recorded an average processing time of 0.01 seconds per segment, in stark contrast to COMSOL’s 7.5 seconds, resulting in a computational speedup of approximately $\mathbf{750\times}$. This demonstrates that our first-stage analytical solver delivers both rapid and precise results when juxtaposed with FEM-based COMSOL simulations.
Fig. 3 illustrates the comparative results of our analytical solution versus COMSOL for segments with and without voids, achieved with $M=5$ in the analytical calculations, which was decided by experimental determination and consistently applied throughout this study.
IV-D Overall Accuracy and Performance Analysis
((a)) Ten-segment interconnect structure. Depicted current densities are the mean values of the current densities used during simulation.
((b)) Node #1 (node with void) in (a).
((c)) Inter-segment junction #2 in (a).
((d)) Inter-segment junction #5 in (a).
((e)) Inter-segment junction #7 in (a).
Figure 4: Illustration of stress variations at junctions of a multi-segment interconnect structure. (a): Example of a ten-segment interconnect. (b), (c), (d), (e): Comparisons of variations estimation between the proposed framework, COMSOL and EMSpice at different junctions.
To assess the efficacy and precision of our method across a variety of general multi-segment interconnect structures, we analyzed a total of 1000 randomly generated configurations, each containing between 50 and 250 segments, with each inter-segment junction connected to at most four segments. These structures were examined to gauge variational EM stress through the use of COMSOL [16], EMSpice [13], and our proposed BPINN-EM-Post, utilizing 30 samples to measure variations in stochastic EM stress. We should notice that other existing works [20, 22] are not applicable for stochastic analysis, while [24] cannot be employed in post-voiding phase, thus these methods are not available for numerical comparison. The findings, detailing accuracy and performance metrics for multi-segment interconnect structures across different segment counts, are showcased in Tab. I. Our method exhibits superior acceleration than prior arts, with error rate of only 1.10% at most, compared with COMSOL and less than 0.2% compared with EMSpice. In addition, Training and Inference time shows nearly linear growth with respect to the number of wires, indicating its scalability to larger interconnect structures.
@@ IV-D Overall Accuracy and Performance Analysis
The observed range of EM stress across these structures spanned from $4.31\times 10^{9}$ Pa to $5.94\times 10^{9}$ Pa. Accuracy was quantified using RMSE defined as $RMSE=(RMSE_{mean}+RMSE_{std})/2$. For dataset comprising 1000 multi-segment interconnect structures, the average RMSE relative to COMSOL stood at $7.07\times 10^{7}$ Pa, equating to an error margin around 0.69%. In comparison, against EMSpice, the average RMSE was recorded at $9.5\times 10^{6}$ Pa, reflecting an error rate about 0.093%.
@@
Fig. 4 demonstrates the variations in EM stress, expressed through both mean and standard deviation, for a ten-segment interconnect structure for simplicity. The physical and electrical characteristics of which are shown in Fig. 4(a). The variations in EM stress at the selected junctions between segments of this interconnect structure are detailed in Fig. 4(b), Fig. 4(c), Fig. 4(d), and Fig. 4(e). In these comparisons, our proposed BPINN-EM-Post is shown to precisely predict the variations in EM stress, aligning closely with the results obtained from COMSOL [16] and EMSpice [13].
@@
In the analysis of stochastic EM stress variances for multi-segment interconnect structures, COMSOL’s average runtime for 1000 multi-segment interconnect structures is approximately 7500 seconds, while EMSpice takes about 2100 seconds in average. For BPINN-EM-Post, variations in stochastic EM stress are estimated by sampling from the BNN via HMC. Prior to this, the analytical solutions must ascertain the optimal atomic flux at the termini of each segment within the structures, leveraging a PINN. Thus, both the sampling (training) period of the PINN and the inference duration of the BNN are accounted for in the runtime evaluation. For the multi-segment interconnects assessed, the average runtime for BPINN-EM-Post is noted to be approximately 31 seconds, significantly accelerating the estimation process by a factor of $240\times$ faster than COMSOL and $67\times$ quicker than EMSpice. This efficiency marks our approach as vastly superior in terms of speed for calculating variational EM stress.
V Conclusion
@@
In this paper, we introduced a novel machine learning-based stochastic analysis framework, termed BPINN-EM-Post, to efficiently estimate variational Electromigration (EM)-induced stress effects. This framework combines closed-form analytical solutions with a PINN to ensure stress continuity and atomic flux conservation during post-void stress calculations. BPINN-EM-Post effectively utilizes analytical solutions for individual wire segments, avoiding the need to model complex initial stress distributions. At the inter-segment level, physics constraints, such as stress continuity and atomic flux conservation, are enforced through the neural network, resulting in a more efficient and accelerated training process by reducing the number of variables in the loss function. The BPINN framework further enhances the approach by providing uncertainty quantification and enabling variational analysis, ensuring robust and efficient modeling of EM stress evolution. Experimental results demonstrate substantial computational improvements, achieving a $240\times$ speedup compared to FEM-based Monte Carlo simulations and a $67\times$ acceleration over FDM-based methods. These findings underscore the potential of the BPINN framework for scalable and accurate variational EM stress analysis.